# BLaST: High Performance Inference and Pretraining using BLock Sparse Transformers

### v1

@@ Abstract.
The energy consumption of large-scale ML models is dominated by data movement - shuffling billions of parameters across memory hierarchies and data centers. Effective sparsification to prune redundant parameters is still challenging: existing methods incur significant accuracy degradation, performance overhead, or both. We introduce (Bl)ock (a)nd (S)parse (T)ransformers (BLaST), a general, robust, and reliable sparsification method applicable to linear layers in all settings. Our method iteratively sparsifies weight matrices into a block sparsity pattern suitable for efficient sparse matrix-matrix (SpMM) multiplication. BLaST achieves up to 95% sparsity in MLP weights with negligible accuracy loss. Our fused, highly optimized Sparse MLP kernel delivers up to 16.7x speedup over dense MLPs across 9 architectures and 8 datasets, resulting in up to 1.6x inference speedup, 1.11x pretraining speedup and up to 3.12x inference memory usage reduction. BLaST enables the next generation of large-scale AI systems by reducing energy use, memory footprint, and latency.
@@ 1. Introduction
Figure 1. BLaST can sparsify MLP weight matrices up to 95% with minimal accuracy loss. Our highly optimized SpMM kernel improves end-to-end inference runtime by up to 1.6x of Llama 3.2-1B. Sparsified matrices reduce the number of GPUs required to process Llama 3.2-405B model by up to 2.9x, leading to significant cost reductions.
Figure 2. A high-level overview of the entire BLaST pipeline. Our method offers the flexibility to either leverage pretrained neural network weights for fine-tuning or initialize the model from scratch for pretraining. BLaST performs a forward and backward pass, updating the weights while leveraging a fused, highly optimized Sparse MLP kernel. Then, for MLP blocks BLaST sparsifies the weight matrices with the blocked prune and grow algorithm. First, weights are divided in blocks of size $b\times b$. The blocks with the highest magnitudes, along with those that exhibit the largest gradient magnitudes, are retained. This procedure repeats until the desired sparsity is reached.
@@
Large Language Models (LLMs) based on the Transformer architecture have driven a dramatic expansion in the adoption of AI across everyday applications. State-of-the-art models (Geminiteam2024, ; Touvron2023a, ; Deepseekai2025, ; Radford2018, ; Radford2019, ; Brown2020, ) rely on large-scale data centers for both pretraining on massive datasets, and subsequent deployment to serve millions of users. While recent algorithmic innovations have reduced model size and computational demands (Ainslie2023, ; Meng2025, ), the trend toward ever-larger parameter counts continues unabated (Shen2025, ). Training such models to high accuracy and serving them with low latency requires substantial investment in hardware infrastructure, making cost reduction a central challenge in democratizing AI access.
These hardware investments are overwhelmingly tailored to Transformer-based architectures (Vaswani2023, ). The associated costs scale with the size of the model’s weight matrices: larger models consume more memory, and require more GPUs for storage and computation. At runtime, the dominant energy expenditure stems from data movement—transferring weights within and between memory and compute units on chip and across systems. Revenue generation, in turn, is tightly coupled to throughput, i.e., the number of tokens produced per second. Model pruning (Mishra2021, ; Evci2021, ; Hoefler2021, ; Frantar2023, ; Vanderschueren2022, ; Hubara2021, ; Zhou2021, ; Ashkboos2024, )—removing redundant parameters—is a principled strategy that simultaneously reduces memory footprint, minimizes data movement, and increases inference throughput. As such, it represents a compelling approach for improving efficiency across training, deployment, and serving of LLMs at scale.
A large body of work has shown that dense neural networks contain significant parameter redundancy, and that selectively pruning these parameters can yield smaller models with negligible accuracy degradation (Hoefler2021, ; Koch2007, ; Schotthofer2022, ; Ceruti2022, ; Ceruti2022a, ; Hu2024, ; Hu2024a, ; Hubara2021, ; Zhou2021, ; Zhou2021a, ; Evci2021, ; Frankle2019, ; Child2019, ; Pan2024, ; Mishra2021, ; Ivanov2023, ; Hu2021, ; Buyukakyuz2024, ; Chen2024, ; castro2023venom, ; Frantar2023, ; Ashkboos2024, ; Chen2021a, ; Chen2021, ; Chen2020, ; You2022, ). For instance, Evci et al. (Evci2021, ) prune ResNet (He2015, ) convolution layers by leveraging weight magnitudes and first-order gradients, achieving up to 90% sparsity in pretraining. Sparse tensor operations also require fewer floating-point operations, reducing both compute and data movement.
@@
Despite active research, developing a general-purpose sparsification method that simultaneously delivers high sparsity, high performance, and minimal accuracy loss remains difficult. The primary challenge lies in hardware constraints: modern accelerators rely on dense, batched memory access patterns and computation. For example, Tensor Cores (Markidis2018, ) are optimized for dense matrix fragments, and features like Tensor Memory Accelerators (TMAs) (Nvidia2022, ) require deep memory pipelines to fully utilize available bandwidth. While high sparsity has been demonstrated through unstructured pruning—removing parameters element-wise—such sparsity patterns are difficult to process efficiently on current hardware. Attempts to enforce structured sparsity (e.g., pruning blocks or entire channels) offer better hardware compatibility but typically fail to exceed 25% sparsity without incurring significant accuracy loss (Ashkboos2024, ). The achievable speedup of N:M sparsity (Mishra2021, ; Hu2024a, ; Hu2024, ; castro2023venom, ) is limited due to the application of sparsity only in contiguous blocks of elements in spite of having hardware support for Nvidia GPUs(Mishra2021, ).
To address these challenges, we introduce BLaST, a block sparse neural network framework designed from the ground up with hardware efficiency in mind. The key idea is to jointly optimize for both algorithmic sparsity and hardware execution by integrating aggressive, accuracy-preserving weight pruning with custom high-performance sparse kernels. BLaST identifies optimal pruning strategies while leveraging a specialized SpMM (Sparse Matrix-Matrix Multiplication) kernel that outperforms even vendor optimized dense (NVIDIA_cuBLAS_2025, ) and sparse (Abdelfattah2024, ; cusparse, ; hipsparse, ) implementations, delivering significant gains. Our kernel accelerates matrix multiplication by up to 16.7x, enhances the MLP layers in Llama-405B by 8.8x, and boosts inference performance by 1.6x—thus enabling the development of trillion-parameter models. Figure 1 shows the key improvements by BLaST to the Llama3.2(Touvron2023a, ) family of models. The high level design of BLaST is presented in Figure 2.
BLaST provides a unified solution for efficient machine learning across pretraining, fine tuning, and post-training compression. We integrate sparsity into the entire model lifecycle while aligning with hardware execution constraints. During pretraining, BLaST applies a custom, hardware-friendly prune-and-grow strategy that incrementally removes redundant weights, reaching up to 80% sparsity with minimal performance overhead—each pruning step simultaneously improves the efficiency of forward and backward passes. In the fine tuning stage, BLaST builds on the sparsified model, and integrates seamlessly with existing training pipelines [^1] https://huggingface.co/docs/transformers/training, achieving up to 95% sparsity while increasing inference throughput. For post-training model compression, BLaST employs knowledge distillation to generate a compact surrogate model from a dense pretrained one, preserving accuracy while reducing memory footprint, and improving forward pass latency the same as in the fine tuning stage.
The custom SpMM kernel developed in BLaST is also available as a standalone sparse linear algebra primitive, designed for use beyond AI workloads. Here, we exploit block sparsity to unlock unprecedented efficiency across a broad range of scientific and industrial applications where sparse computation is increasingly critical. BLaST’s blocked SpMM (BSpMM) sets a new benchmark, outperforming all existing SpMM libraries—both open-source and commercial. Remarkably, for favorable input structures, BSpMM can even surpass CUTLASS in dense matrix-matrix multiplication performance. On average, as little as 50% sparsity is sufficient to outperform vendor-optimized GEMM libraries, and at 95% sparsity—commonly achieved by BLaST—BSpMM delivers up to 16.7x speedup over the best existing dense and sparse alternatives.
@@
  1. (1)
    We introduce BLaST, a block sparse neural network framework designed from the ground up for hardware efficiency. BLaST provides a unified, end-to-end solution for efficient machine learning, covering the full model lifecycle—including pretraining, fine tuning, and post-training compression—while maintaining high accuracy under extreme sparsity.
  2. (2)
    We develop a high-performance, scalable implementation of BLaST, including a custom block-sparse SpMM kernel optimized for modern GPU architectures that significantly outperforms both dense and existing sparse baselines.
  3. (3)
    We conduct an extensive empirical evaluation across 15 model architectures and 9 datasets, demonstrating that BLaST achieves up to 90–95% sparsity with no accuracy degradation. Our results show up to 16.7x speedup in kernel performance and up to 1.6x improvement in end-to-end training and inference throughput. BLaST achieves up to 3.12x reduction in inference memory footprint, resulting in 2.9x reduction in GPU setup and operating costs.
@@ 2. Background
We now review the key components of the Transformer-based models. We also analyze them from the scope of sparsification strategies and lay foundations for the BLaST approach.
@@ 2.1. Transformer-based ML models
The Transformer architecture (Vaswani2023, ) has emerged as the foundational building block of state-of-the-art large language models (Touvron2023a, ; Deepseekai2025, ; Geminiteam2024, ),vision models (dosovitskiy2020image, ), models for science (jumper2021highly, ; sonderby2020metnet, ), and multimodal models
 (radford2021learning, ; ramesh2021zero, ). Each layer of the Transformer architecture comprises two main components: Attention and MLP, plus normalization layers (Figure 2).
Attention. While there are many variants of the Attention module: multi-head attention (Vaswani2023, ), multi-query attention (shazeer2019fast, ), grouped-query attention (ainslie2023gqa, ) and multi-latent attention (liu2024deepseek, ), its underlying compute pattern is the multiplication of transient tensors $Q$, $K$, and $V$ that are computed based on the activation tensor $X$ and trainable weight matrices. The Attention layer captures the relations between tokens - the output state of the token after the attention module depends on the values of other tokens.
The MLP layer, on the other hand, multiplies the activation tensor $X$ directly with weight matrices. Depending on the architecture, typically there are two (radford2019language, ) or three (Touvron2023a, ) weight matrices. Arguably, MoE (mixture-of-experts) (shazeer2017outrageously, ) can be viewed as a functional equivariant of MLP: the essence being that each token is processed independently - the state of the token after MLP depends only on its previous state and the weight matrices, but does not depend on other tokens. For the Llama-family models, the MLP operator is defined as:
@@
where $SiLU$ represents the Sigmoid Linear Unit activation function, $X$ represents the input to the MLP layer, and $W_{1}$, $W_{2}$, and $W_{3}$ are the weight matrices. The MLP layer accepts a tensor $X,X\in\mathbb{R}^{n\times s\times e}$ of size $n\times s\times e$, where $n,s,e$ are batch size, sequence length, and embedding size, respectively. Then the model performs a series of expansions and contractions on this input $X$ to generate the output $Y$ as shown in Equation 1.
@@ 2.2. Model Sparsification
Related works on sparsification in deep learning suggest that no single method is universally applicable across all scenarios (hoefler2021sparsity, ). Various techniques have been explored—from unstructured weight pruning (Frankle2019, ; prasanna2020bert, ; gordon2020compressing, ; Chen2020, ; guo1909reweighted, ; sanh2020movement, ), which eliminates redundant individual connections, to structured approaches that remove entire neurons (polyak2015channel, ; hu1607network, ; pan2016dropneuron, ; lee2018snip, ), filters (luo2020autopruner, ; luo2017thinet, ; he2019filter, ; you2019gate, ), or methods that sparsify attention (michel2019sixteen, ; xu2025xattention, ; voita2019analyzing, ; mccarley2020structured, ; yun2020n, ; parmar2018image, ; guo2019star, ; child2019generating, ; beltagy2020longformer, ; zaheer2020big, ; li2020sac, ; zhao2019explicit, ). Recently, XAttention (xu2025xattention, ) has proposed block sparse attention with antidiagonal scoring, and NSA (yuan2025native, ) introduces a natively trainable sparse attention mechanism – both for long-context Transformers.
However, empirical studies indicate that either sparsifying MLP weights or adopting a Mixture-of-Experts (MoE) tends to be more robust, leading to substantial reductions in memory footprint and improvements in parallelization (shazeer2017outrageously, ; lepikhin2020gshard, ; fedus2022switch, ; huang2024toward, ). In contrast, approaches focused on sparsifying attention components mainly reduce the size of transient storage during computation without significantly lowering the overall parameter count. These observations underscore the need for flexible, multi-strategy sparsification approaches that balance efficiency gains with the preservation of network trainability and generalization.
In this work, we propose a general solution that is applicable across the entire range of current and future Transformer architectures. Our proposed solution remains accurate as sparsity in the MLP layers increases, thereby ensuring scalable performance for the future of AI.
@@ 3. BLaST - (BL)ock (S)parse (T)ransformers
We now proceed to describe BLaST — a novel method designed to improve both inference and pretraining execution time, while also reducing the memory footprint. Section 3.2 introduces the algorithm to obtain the neural networks in the desired block sparse pattern. Section 3.3 introduces our optimized BSpMM kernel.
@@ 3.1. Overview
Figure 2 illustrates the high-level pipeline of BLaST for pretraining and finetuning. At the start of training, if pretrained weights are available, they are used to initialize the Transformer-based architecture. Next, BLaST performs a forward and backward pass using our optimized BSpMM kernel (Section 3.3), updating the parameters. Then, the blocked prune and grow algorithm (Section 3.2) is applied in the MLP blocks, adaptively pruning and regrowing weights. This process repeats until the desired sparsity level is reached.
3.2. Blocked Prune and Grow
The “early bird” lottery ticket hypothesis (Chen2021, ; You2022, ) proposes that certain weights are more important than others, and that they can be selected during pretraining without significantly affecting the model’s accuracy. According to this hypothesis, winning weights are chosen by pruning the weights with the smallest magnitudes (Frankle2019, ; Chen2021a, ; Chen2021, ; You2022, ). RigL (Evci2021, ; Evci2022, ) demonstrated that the pruned model’s accuracy can be further improved by incorporating gradient flow information to select the winning weights. This improvement is achieved by applying the same pruning function, $S()$, to both the weights and their corresponding gradients. In the final step, a weight is retained if it meets one of two conditions: it is either directly preserved by the function $S()$ or regrown based on gradient information.
Using RigL on CNNs (Evci2021, ; Evci2022, ) results in fine-grained unstructured sparsity of the weight matrices. Although this results in a reduced number of floating point operations during the forward and backward propagation, it does not translate into improved performance due to memory bandwidth limitations of modern GPUs(Abdelfattah2024, ; naumov2010cusparse, ; hipsparse, ). As a result, we propose to prune the weight matrices of the MLP layer of LLMs using block sparse matrices as shown in the bottom part of Figure 2.
Our blocked prune and grow algorithm gradually reduces the number of parameters of the weight matrix $W$ until a sparsity $s_{max}$ has been achieved over $m$ iterations. It can be implemented as a simple extension to a forward and backward training loop. Listing LABEL:lst:loop-prune-and-grow shows the blocked prune and grow procedure with a regular training loop as Python code. The training loop and optimizer are executed in forward_and_backward_step(). The masks are updated once every $step\_size$ iterations. The masks for each weight matrix are stored in boolean tensors and applied to the weight matrices in the prune_weights() step.
@@ 3.2. Blocked Prune and Grow
4 generate_masks()
5 prune_weights()
Listing 1: Training loop of our blocked prune and grow algorithm. The actual grow and prune procedure is implemented in the function generate_masks().
Figure 2 illustrates the prune and grow procedure implemented by the function generate_masks() for a weight matrix $W_{i}$ and its corresponding gradient $G_{i}$ for a single iteration. First, the matrices are interpreted as blocks of size $b\times b$, after which the pruning function $S()$ is applied to both $W_{i}$ and $G_{i}$. The function $S()$ computes the Frobenius norm of each block and drops the blocks with the least norm in order to maintain a sparsity $s_{i}$ as given by Equation 2. The pruned blocks which will be zeroed are represented by the white blocks in Figure 2. The third difference step computes the set difference between the non-zero block indices in the $S(W_{i})$ and $S(G_{i})$ matrices, denoted by the red blocks in matrix $D_{i}$. These blocks are regrown back into the sparse weight matrix $S(W_{i})$ in the fourth step, which results in the newly pruned weight matrix $W_{new}$. The regrown blocks added to $W_{new}$ are initially set to zero and then updated during subsequent optimization steps in training. The initialization to zero is done so that the regrown weights do not contribute to the transformation applied by the matrix. Note that in our actual implementation we use boolean masks throughout the pruning and growing process to save time and memory. The $W_{new}$ matrix is ultimately generated in the prune_weights() function.
The weight sparsity $s_{i}$ at iteration $i$ is determined by Equation 2 (Zhu2017, ).
@@
where $s_{init}$ represents the initial sparsity, which determines the sparsity at the beginning of the training, $m$ is the maximum number of training iterations, and $d$ is a decay term that controls how fast the maximum sparsity $s_{max}$ is reached.
Unlike pruning algorithms that use straight-through estimators (Hu2024a, ; Hubara2021, ; Bengio2013, ), our method prunes the weight matrices directly. Consequently, the same pruned weight matrix is applied during both forward and backward propagation. During training, the dense weight and gradient matrices are kept intact, while the pruned weight $W_{new}$ is stored in the blocked compressed sparse column (BCSC) format [^2] https://pytorch.org/docs/stable/sparse.html#sparse-bsc-tensor for forward and backward propagation. Although this leads to higher memory consumption during training (which gradually decreases as the sparsity $s_{i}$ increases), it ultimately yields improved speedup and reduced inference memory usage, as shown in Section 5.

### v2

@@ Abstract.
The energy consumption of large-scale ML models is dominated by data movement, shuffling billions of parameters across memory hierarchies and data centers. Sparsification offers a principled way to mitigate these costs by pruning redundant weights and activations, thereby reducing data movement. Effective sparsification to prune redundant parameters is still challenging: existing methods incur significant accuracy degradation, performance overhead, or both. We introduce (Bl)ock (a)nd (S)parse (T)ransformers (BLaST), a general, robust, and reliable method for sparsification, applicable to linear layers in all settings. Our method iteratively sparsifies weight matrices into a block sparsity pattern suitable for efficient sparse matrix-matrix (SpMM) multiplication. BLaST achieves up to 95% sparsity in MLP weights with negligible accuracy loss (majority <<2.25%). We show a 2.2x inference speedup for Llama 3.2 with 16 GPUs, and up to 4.45x reduction in inference memory footprint resulting in a 2.9x reduction in GPU setup and operating costs.
@@ 1. Introduction
Large Language Models (LLMs) based on the Transformer architecture have driven a dramatic expansion in the adoption of AI across everyday applications. State-of-the-art models (Geminiteam2024, ; Touvron2023a, ; Deepseekai2025, ; Radford2018, ; Radford2019, ; Brown2020, ) rely on large-scale data centers for both pretraining on massive datasets, and subsequent deployment to serve millions of users. While recent algorithmic innovations have reduced model size and computational demands (Ainslie2023, ; Meng2025, ), the trend toward ever-larger parameter counts continues unabated (Shen2025, ). Training such models to high accuracy and serving them with low latency requires substantial investment in hardware infrastructure, making cost reduction a central challenge. These hardware investments are overwhelmingly tailored to Transformer-based architectures (Vaswani2023, ). The associated costs scale with the size of the model’s weight matrices: larger models consume more memory, and require more GPUs for storage and computation.
[…]
Figure 2. Overview of BLaST. Training or fine-tuning alternates dense updates with a pluggable block-pruning stage (instantiated with Blocked Prune-and-Grow (Section 3.2) and oLLM (Section 3.3)) while using a fused, high-performance sparse MLP kernel (Section 3.4) until the target sparsity is reached.
@@
Model pruning (Hoefler2021, ; Koch2007, ; Schotthofer2022, ; Ceruti2022, ; Ceruti2022a, ; Hu2024, ; Hu2024a, ; Hubara2021, ; Zhou2021, ; Zhou2021a, ; Evci2021, ; Frankle2019, ; Child2019, ; Pan2024, ; Mishra2021, ; Ivanov2023, ; Hu2021, ; Buyukakyuz2024, ; Chen2024, ; castro2023venom, ; Frantar2023, ; Ashkboos2024, ; Chen2021a, ; Chen2021, ; Chen2020, ; You2022, ), i.e. removing redundant parameters, is a principled strategy that simultaneously reduces memory footprint, minimizes data movement, and increases inference throughput. As such, it represents a compelling approach for improving efficiency across training, deployment, and serving of LLMs at scale. Despite active research, developing a general-purpose sparsification method that simultaneously delivers high sparsity, high performance, and minimal accuracy loss remains difficult. The primary challenge lies in hardware constraints: modern accelerators rely on dense, batched memory access patterns and computation. Unstructured pruning can reach high sparsity by removing individual weights, but irregular patterns are hard to accelerate on current hardware (Frankle2019, ; prasanna2020bert, ; gordon2020compressing, ; Chen2020, ; guo1909reweighted, ; sanh2020movement, ). Attempts to enforce structured sparsity (e.g., pruning blocks or entire channels) offer better hardware compatibility but typically fail to exceed sufficient sparsity for speedups without incurring significant accuracy loss (Ashkboos2024, ). Even with NVIDIA GPU support (Mishra2021, ), N:M sparsity prunes only within fixed-size contiguous groups, which limits speedup (Hu2024a, ; Hu2024, ; castro2023venom, ).
[…]
@@
  • Framework: We introduce BLaST, a block sparse neural network framework designed from the ground up for hardware efficiency. BLaST provides a unified, end-to-end solution for efficient machine learning, covering the full model lifecycle;
  • Kernel: We develop a high-performance, scalable implementation of a custom block-sparse SpMM kernel optimized for modern GPU architectures;
  • Evaluation: We conduct an extensive empirical evaluation across 8 datasets and 11 model architectures, outperforming state-of-the-art kernels while preserving accuracy across tasks.
@@ 2. Background
We now review the key components of the Transformer-based models. We also analyze them from the scope of sparsification strategies, and lay the foundations for BLaST.
@@ 2.1. Transformer-based ML models
The Transformer architecture (Vaswani2023, ) has emerged as the foundational building block of state-of-the-art large language models (Touvron2023a, ; Deepseekai2025, ; Geminiteam2024, ), vision models (dosovitskiy2020image, ), models for science (jumper2021highly, ; sonderby2020metnet, ), and multimodal models (radford2021learning, ; ramesh2021zero, ). Each block of the Transformer architecture comprises two main components: Attention and MLP, along with normalization layers (Figure 2).
Attention. The attention layer (shazeer2019fast, ; Vaswani2023, ; ainslie2023gqa, ; liu2024deepseek, ) captures relationships between tokens, where each token’s output representation is computed based on the values of all other tokens in the sequence.
The MLP layer, on the other hand, multiplies the activation tensor $X$ directly with weight matrices. Depending on the architecture, typically there are two (radford2019language, ) or three (Touvron2023a, ) weight matrices. A Mixture-of-Experts (MoE) (shazeer2017outrageously, ) can be viewed as functionally analogous to a standard MLP layer, as both operate on each token independently. In either architecture, the output state of a token is determined solely by its input state and the network’s weights, without influence from other tokens in the sequence. For the Llama3 family models, the MLP operator is defined as:
@@
where $SiLU$ represents the Sigmoid Linear Unit activation function, $X$ represents the input to the MLP layer, and $W_{1}$, $W_{2}$, and $W_{3}$ are the weight matrices. The MLP layer accepts a tensor $X\in\mathbb{R}^{n\times s\times e}$ of size $n\times s\times e$, where $n,s,e$ are batch size, sequence length, and embedding size, respectively. Then the model performs a series of expansions and contractions on this input $X$ to generate the output $Y$ as shown in Equation 1.
@@ 2.2. Model Sparsification
Related works on sparsification in deep learning suggest that no single method is universally applicable across all scenarios (hoefler2021sparsity, ). Various techniques have been explored, from unstructured weight pruning (Frankle2019, ; prasanna2020bert, ; gordon2020compressing, ; Chen2020, ; guo1909reweighted, ; sanh2020movement, ), which eliminates redundant individual connections, to structured approaches that remove entire neurons (polyak2015channel, ; hu1607network, ; pan2016dropneuron, ; lee2018snip, ), filters (luo2020autopruner, ; luo2017thinet, ; he2019filter, ; you2019gate, ), or methods that sparsify attention (michel2019sixteen, ; xu2025xattention, ; voita2019analyzing, ; mccarley2020structured, ; yun2020n, ; parmar2018image, ; guo2019star, ; child2019generating, ; beltagy2020longformer, ; zaheer2020big, ; li2020sac, ; zhao2019explicit, ). Recently, XAttention (xu2025xattention, ) has proposed block sparse attention with antidiagonal scoring, and NSA (yuan2025native, ) introduces a natively trainable sparse attention mechanism, both for long-context Transformers.
However, empirical studies indicate that sparsifying MLP weights tends to be more robust, leading to substantial reductions in memory footprint and improvements in parallelization (shazeer2017outrageously, ; lepikhin2020gshard, ; fedus2022switch, ; huang2024toward, ). In contrast, approaches focused on sparsifying attention components mainly reduce the size of transient storage during computation without significantly lowering the overall parameter count. These observations underscore the need for flexible, multi-strategy sparsification approaches that balance efficiency gains with the preservation of network trainability and generalization.
In this work, we propose a general solution that is applicable across the entire range of current and future Transformer architectures. Our proposed solution remains accurate as sparsity in the MLP layers increases, thereby ensuring scalable performance for the future of AI. Although our experiments focus on MLP sparsification, the framework accommodates sparsity in any linear layer.
@@ 3.1. Overview
We now describe BLaST, our novel framework for improving the pre-training and inference execution time, while also reducing memory footprint. The sparsification algorithms are described in Section 3.2 and Section 3.3, and the BSpMM kernel in Section 3.4.
3.2. Blocked Prune and Grow (P&G)
Magnitude pruning can expose early subnets during pretraining (Frankle2019, ; Chen2021a, ; Chen2021, ; You2022, ). RigL improves accuracy by regrowing connections using gradients (Evci2021, ; Evci2022, ). These methods yield unstructured sparsity that is hard to accelerate on GPUs due to bandwidth limits (Abdelfattah2024, ; naumov2010cusparse, ; hipsparse, ). We adapt prune and grow to block-sparse weights in Transformer MLPs, as shown in Figure 2, and refer to it as P&G.
The P&G algorithm gradually reduces the number of parameters of the weight matrix $W$ until a sparsity $s_{max}$ has been achieved over $m$ iterations. Listing LABEL:lst:loop-prune-and-grow shows the algorithm with a regular training loop as Python code. The training loop and optimizer are executed in forward_and_backward_step(). The masks are updated once every step_sizestep\_size iterations. The masks for each weight matrix are stored in boolean tensors and applied to the weight matrices in the prune_weights() step.
@@ 3.2. Blocked Prune and Grow (P&G)
4 masks = generate_masks()
5 prune_weights(masks)
Listing 1: Training loop of the blocked prune and grow algorithm. The actual grow and prune procedure is implemented in the function generate_masks().
Figure 2 shows one iteration of generate_masks() on weight $W_{i}$ and gradient $G_{i}$. We partition $W_{i}$ and $G_{i}$ into $b\times b$ blocks and apply $S(\cdot)$. Each block is scored by its Frobenius norm. We drop the lowest scoring blocks until we reach sparsity $s_{i}$, as given in Equation 2. Zeroed blocks are shown in white in the figure. Next, we take the set difference of the nonzero block indices between $S(W_{i})$ and $S(G_{i})$. The resulting indices, shown in red in matrix $D_{i}$, are regrown into $S(W_{i})$, which yields the updated weight $W_{\text{new}}$. Newly regrown blocks are initialized to zero so that they do not affect the current transform and are updated during subsequent optimization. In practice, we keep boolean masks for pruning and growth and form $W_{\text{new}}$ in prune_weights().
The weight sparsity $s_{i}\in[0,1)$ at iteration $i$ is determined by Equation 2 (Zhu2017, ).
@@
where $s_{init}$ represents the initial sparsity, which determines the sparsity at the beginning of the training, $m$ is the maximum number of training iterations, and $d$ is a decay term that controls how fast the maximum sparsity $s_{max}$ is reached. Cubic scaling provides for a smooth and gradual increase in sparsity.
Unlike pruning algorithms that use straight-through estimators (Hu2024a, ; Hubara2021, ; Bengio2013, ), our method prunes the weight matrices directly. Consequently, the same pruned weight matrix is applied during both forward and backward propagation. During training, the dense weight and gradient matrices are kept intact, while the pruned weight $W_{new}$ is stored in the blocked compressed sparse column (BCSC) format (im2000optimizing, ) for forward and backward propagation. Although this leads to higher memory consumption during training (which gradually decreases as the sparsity $s_{i}$ increases), it ultimately yields improved speedup and reduced inference memory usage, as shown in Section 5.
3.3. oLLM
We propose a method called “oLLM” that uses second-order gradient information for pruning weight blocks within layers, inspired by the optimal brain surgeon method (Hassibi1993, ). The OBS based sparsification process consists of four steps: first, estimate the model’s curvature; second, assign sensitivity scores to each weight based on this curvature; third, prune weights with the lowest sensitivity; and fourth, update the remaining nonzero weights to preserve the curvature estimated by the second-order gradient.
The core principle of OBS is to estimate the sensitivity $\mathcal{L}_{k}$ of the $k$-th weight using the diagonal element $H^{-1}_{kk}$ of the inverse Hessian matrix $H^{-1}$, as shown in Equation 3. Weights with lower sensitivity can be pruned with minimal accuracy loss.
(3) $\mathcal{L}_{k}=\frac{1}{2}\frac{W_{k}^{2}}{[H^{-1}]_{kk}}$
The remaining weights are then updated as shown in Equation 4, where $e_{k}$ is a basis matrix for selecting the corresponding column of $H^{-1}$. The size of the Hessian matrix is a square of the number of weights, which makes it intractable to store and invert in a finite amount of time. To overcome this limitation, the diagonal inverse Fisher approximation(Gupta2018, ; Kurtic2022, ; Grosse2016, ; Vanderouderaa2024, ) can be used for approximating the inverse Hessian $H^{-1}$ to save space and computation time.
(4) $\Delta W=-\frac{W_{k}}{[H^{-1}]_{kk}}H^{-1}e_{k}$
Figure 3. The blocked oBERT accumulates weight gradients $G_{i}$ over many forward and backward iterations to generate the sensitivity scores $S_{i}$ using the Fisher inverse matrix. $S_{i}$ is then used for generating a blocked mask $M_{i}$. The blocked weight matrix $W_{new}$ is generated with an element-wise multiplication of the weight matrix $W_{i}$ with $M_{i}$.
Second order pruning methods such as oBERT (Kurtic2022, ) assume that the first order gradient is already at the minimum. On the contrary, methods such as LPViT (xu2024lpvit, ) incorporate the first order gradient information into the sensitivity score calculation and show better accuracy. Equation 5 shows the calculation of the sensitivity $\mathcal{R}_{k}$ using the first order gradient for the $k^{th}$ weight in weight matrix $W$. $\bar{G}$ is the accumulated first order gradient over $g$ iterations as shown in Figure 3. Equation 6 then computes the total sensitivity $S_{k}$ using $\mathcal{L}_{k}$ and $\mathcal{R}_{k}$.
(5) $\mathcal{R}_{k}=|W_{k}\cdot\bar{G}_{k}|$
(6) $\mathcal{S}_{k}=\mathcal{L}_{k}+\mathcal{R}_{k}$
The weight sensitivity $S_{k}$ of each weight is combined to form the sensitivity matrix $S_{i}$ corresponding to each weight in the weight matrix $W_{i}$ as shown in Figure 3. $S_{i}$ is then split into blocks of size $b\times b$, and the sensitivity score of the block is obtained with 2D average pooling (zafar2022comparison, ). The weights corresponding to blocks with the least block sensitivity are then pruned to obtain the block sparse weight matrix $W_{new}$.